slipknot3066:

Evaluate -2[7 - 3(-8 + 5)]

Hero:

Apply PEMDAS Evaluate everything in parentheses first beginning with the inner most parentheses: So first perform the operation -8 + 5 what do you get?

slipknot3066:

-3

Hero:

Okay, so now you have -2[7 - 3(-3)] or to make it more clear, you're left with \(-2[7 - 3 \times -3]\) So you have a product and a multiplication within the grouping symbols. According to PEMDAS, which operation do you think should be performed first: Multiplication or Subtraction?

slipknot3066:

Multiplication

Hero:

Correct and what do we multiply?

slipknot3066:

-2 x 7-3

Hero:

Actually, we need to deal with what is within the grouping symbols. The only thing you need to focus on is this expression for now: \(7 - 3 \times -3\)

Hero:

Knowing this, what operation do we perform?

slipknot3066:

7-3 x -3 = 16 x -2 = -32 witch is one of the answers.

Hero:

Here is where we are so far: $$\displaystyle \begin{array}{ l c } =-2[ 7\ -\ 3( -8\ +\ 5)] & \mathbf{Given}\\ =-2[ 7\ -\ 3( -3)] & \mathbf{Add\ -8\ and\ 5}\\ =-2[ 7\ -\ 3\ \times \ -3] & \mathbf{-3( -3) \ is\ the\ same\ as\ -3\ \times \ -3} \ \\ & \\ & \end{array}$$

Hero:

All you have to do is just perform the next operation which is \(-3 \times -3\) For now, don't worry about what the answer choices are just yet. That's a distraction for you atm.

slipknot3066:

-3 x -3 is just 9

Hero:

Exactly. So now we have -2[7 + 9]

Hero:

What do we do next?

Hero:

Keep in mind what is inside the grouping symbols always has priority.

slipknot3066:

7+9

Hero:

Yes what is 7 + 9?

slipknot3066:

16 -.-

Hero:

What operation is that?

slipknot3066:

times -2

Hero:

Whoops my bad

Hero:

We have -2[16] sorry.

Hero:

And what is the final answer now?

slipknot3066:

Again -32

Hero:

Correct. But this time we have done the proper steps to get to the -32

slipknot3066:

Thank you

Hero:

You're welcome. Just remember to do the inside operations before you do the outside operations.